# The Developments and Challenges towards Dexterous and Embodied Robotic Manipulation: A Survey

### v1

@@ VI Open Challenges and New Trends
Considering the above challenges, achieving precise capture of the human hand’s 20+ DoFs remains an extremely challenging problem. Thus to acquire high quality datasets for multi-fingered dexterous hand, a promising trend is to build a high-DoF force-tactile agile teleoperation robot system with tight "Human-Robot-Environment" coupling. Such a system can help bridge both the Sim-to-Real and Human-to-Robot gaps in multi-fingered dexterous manipulation. In addition, except for visual modal, integrating other modal sensing (e.g., force, tactile) can enrich contact information, thereby improving manipulation performance.
@@
Second, the robustness of the static learning framework is weak. 1) Current manipulation skill learning frameworks that based on RL/IL can only statically extract manipulation skills from existing datasets, unable to dynamically adjust the learned skills. 2) Integrating humans effectively into the learning loop also remains a challenge. In existing methods, human supervision and guidance are weakly coupled and could not be applied to the learning processes timely, which makes it difficult to accurately reflect humans’ satisfaction with the skill learning. 3) Owning to the mechanisms of RL/IL, when the environments or tasks are beyond the dataset’s coverage, their performance would degrade significantly. Thus to overcome the limitations of static learning frameworks, a promising trend is to incorporate humans into the learning loop, enabling robots to acquire more robust manipulation skills. However, the challenge is how can we reflect the humans’ intensions into the learning loop. To achieve this, integrating humans’ feedback into the learning loop would be a feasible approach. Human’s feedback can be quantified through physiological properties (e.g., body stiffness, eye tracking) that can reflect humans’ intensions. Among all physiological properties, body stiffness is particularly suitable due to its convenience of acquisition and rich representation of human muscle activity. Leveraging stiffness-based intensions recognition, robots can receive feedback in time, which enhances skills learning and facilitates skills transfer from human to robots.
Third, the generalization ability of end-to-end learning framework is insufficient. 1) Current robot manipulation skill learning predominantly adopts an end-to-end learning framework. However, this learning framework suffers from issues such as high task-coupling and lack of explainability. These issues make the end-to-end learning framework struggle to adapt to long-horizontal sequential tasks scenarios. 2) In recent years, while some research has begun exploring skill composition for long-horizontal sequential tasks, it is still limited to simple robotic arm and two-fingered gripper configurations. Research on compositional skills for multi-fingered dexterous hands is still in its early days. Scientific research[88] shows that humans do not follow an end-to-end framework in learning to perform long-horizontal sequential tasks. Instead, when facing long-horizontal sequential tasks, humans naturally perform the task decomposition and then execute the primitive skills automatically and unconsciously. Inspired by this mechanism, a novel robotic manipulation learning framework, Helix framework, is proposed. This framework combines System 1 (Fast Reactive Control) and System 2 (Slower Deliberative Reasoning). However, despite the above framework is proposed, the mechanism of this framework, and how to implement this framework are still underdeveloped. Thus a promising trend is enabling robots to simulate the combined mechanism of humans’ cognitive systems. Building a hybrid intelligent system could be a feasible way for robotic dexterous manipulation learning in long-horizontal sequential tasks scenarios. The above system integrate an upper-level decision-making module and utilize multi-modal sensing as external perceptions. These capabilities would promote the decision-making process.

### v2

@@ VI Open Challenges and New Trends
Considering the above challenges, achieving precise capture of multi-modal datasets for multi-fingered dexterous robotic hands remains a non-trivial task. A promising trend for data collection is to develop a high-DoF haptic agile teleoperation robot system with tight “Human-Robot-Environment" coupling. Some studies have already begun to explore this potential solution. In [98], a teleoperation system with haptic feedback is developed. In this system, a custom-designed 10-DoFs robotic hand with tactile sensing is teleoperated to perform tasks. During task execution, pressure sensors on the fingertips of the robotic hand collect contact force data, while haptic feedback actuators located on the operator’s fingertips display the contact forces. However, the robotic hand used has only 10 DoFs, fewer than a truly dexterous hand (which typically has more than 20 DoFs, e.g., the Shadow Hand). Moreover, in [36], a high-DoF teleoperation system is built to gather training data for robot skill learning. During data collection, tactile information, joint angle states, and end-effector states are recorded. Such a system can address both the Sim-to-Real and Human-to-Robot gaps in multi-fingered dexterous manipulation. In addition, except for visual modal, integrating other modal sensing (e.g., force, tactile) can enrich contact information, thereby improving manipulation performance. Therefore, we believe the haptic agile teleoperation systems are highly promising for building high-quality datasets for multi-fingered dexterous hands.
@@
Second, the robustness of the static learning framework is weak. 1) Current manipulation skill learning frameworks that based on RL/IL can only statically acquire or extract manipulation skills through trial-and-error exploration or from existing datasets, unable to dynamically adjust the learned skills. 2) Integrating humans effectively into the learning loop also remains a challenge. In existing methods, human supervision and guidance are weakly coupled, which makes it difficult to accurately reflect humans’ satisfaction with the skill learning. 3) Owning to the mechanisms of RL/IL, when the environments or tasks are beyond the dataset’s coverage, their performance would degrade significantly. Thus to overcome the limitations of static learning frameworks, a promising trend is to incorporate humans tightly into the learning loop, enabling robots to acquire more robust manipulation skills. As mentioned above, using human’s guidance to promote learning has already been studied and emerged as a trend. In [99], human’s guidance is indicated with discontinuous binary signals. These binary signals are generated via keyboard key pressing. For instance, if the operator satisfies with the current performance, he or she will give a positive signal. On the contrary, if the performance is not desired, a negative signal will be presented. In [100], a human is asked to provide preferences between different skill sequences. Each preference is also indicated by a binary signal, where the preferred sequence is labeled as positive and the other as negative. These preferences are then used to train a reward model, enabling it to assign higher reward values to the human-preferred sequence. The trained reward model substitutes the hand-crafted reward function in a RL framework. Consequently, the RL framework learns to automatically select the skill sequence that a human is most likely to prefer.
Although using discontinuous binary signals as human’s feedback has achieved great successes. They are weakly coupled and have the natural disadvantage of not being able to express the extent of human satisfactoriness. On the other hand, due to the convenience of acquisition and rich representation of human muscle activity. The stiffness of human arm endpoint has already used in many applications [101, 102]. However, using stiffness to indicate human’s guidance and preference in the learning process is still understudied. Thus we believe that using stiffness to represent human’s guidance and preference to enhance skills learning and facilitate skills transfer from human to robots is a worthwhile research direction and a promising trend in the learning processes.
Third, the generalization ability of end-to-end learning framework is insufficient. Many robot manipulation skill learning predominantly adopts an end-to-end learning framework. However, this learning framework suffers from issues such as high task-coupling and lack of explainability. These issues make the end-to-end learning framework struggle to adapt to long-horizontal sequential tasks scenarios.
Scientific researches [103, 104] show that humans do not follow an end-to-end framework in learning to perform long-horizontal sequential tasks. Instead, when facing long-horizontal sequential tasks, humans naturally perform the task decomposition and then execute the primitive skills automatically and unconsciously. Inspired by this mechanism, a novel robotic manipulation learning framework, Helix framework, is proposed. This framework combines System 1 (Fast Reactive Control) and System 2 (Slower Deliberative Reasoning). Previous studies have already explored hierarchical frameworks which are similar to the Helix framework to address long-horizon tasks. Some studies used two-level hierarchical frameworks. For instance, in [105], the proposed framework integrates a low-level primitive library of continuously parameterized skills with a high-level policy. The low-level library contains essential skills for assembly tasks, such as grasping and inserting, implemented using both RL and model-based controllers. The high-level policy, learned via imitation learning from a small set of demonstrations, selects appropriate primitive skills and instantiates them with continuous input parameters. On the other hand, some studies used three-level hierarchical frameworks [106, 107]. It is noteworthy that different studies may adopt different definitions of the functionalities of these levels. In [106], the lowest level conducts primitive learning, the middle level uses the learned primitives to generate robotic skills, and the highest level adopts a Vision-Language Model (VLM) to decompose long-horizontal tasks into separate robotic skills, meanwhile, these robotic skills are planed and formed in to an action sequence for execution. In [107], at the highest level, a Vision-Language-Planning (VLP) agent performs task decomposition and planning. This agent integrates semantic instructions, visual observations, and spatial intelligence to generate a sequential plan of sub-tasks. Subsequently, at the middle level, a semantic abstraction module maps each sub-task into a generalized atomic skill definition. Finally, at the lowest level, a Vision-Language-Action (VLA) model is fine-tuned to execute these atomic skills, thereby completing the original task. These hierarchical frameworks have demonstrate well performance in the long-horizontal tasks.
However, above studies encountered a common limitation that the end-effectors are mainly two-fingered grippers. Neither dexterous high-DoF robotic hands nor haptic information are incorporated. Research on compositional skills for multi-fingered dexterous hands is still in its early days. Thus we believe that integrating the task reasoning and fast control capabilities of hierarchical frameworks with high-DoF haptic hands is promising. Such integration enables dexterous object manipulation and could significantly improve overall system performance. Therefore, we consider the incorporation of multi-fingered dexterous hands and multi-modal sensing into hierarchical learning frameworks to be a highly promising trend for future research.